OpenStudy (anonymous):

n astronaut having mass 320 kg with equipment included is attempting an untethered space walk. The astronaut is initially at rest with respect to the nearby spacecraft, but discharges 30.0 g of nitrogen gas at 900.0 m/s from a portable propulsion unit. How fast is the astronaut moving after doing that? 0.36 m/s, 0.084 m/s , 27 m/s, or 84 m/s ?? :/

OpenStudy (michele_laino):

here we have to apply the momentum conservation law

OpenStudy (anonymous):

ok!

OpenStudy (michele_laino):

so we can write this: \[P = 0.03 \times 900 = ...\] is the momentum of the gas discharged

OpenStudy (michele_laino):

oops..discharged gas

OpenStudy (michele_laino):

how much is P?

OpenStudy (anonymous):

errr 27?

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

the momentum gained by our astronaut is: \[P' = 320 \times V\] where V is his velocity

OpenStudy (michele_laino):

so we can write: \[27 = 320 \times V\] what is V?

OpenStudy (anonymous):

0.984375? so our solution is 0.084 m/s ?

OpenStudy (michele_laino):

\[V = \frac{{27}}{{320}} = {\text{0}}{\text{.084375}}\]

OpenStudy (michele_laino):

so V=0.084

OpenStudy (anonymous):

so is that our final solution? :O

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

yay!! thank you!!

OpenStudy (michele_laino):

thank you!! :)